OpenStudy (anonymous):

Find all values of 'k' for which kx+4y+z=0 4x+ky+2z=0 2x+2y+z=0 have a common line of intersection . Please explain in detail .

OpenStudy (anonymous):

solve q3 for z then put the value in eq 1

OpenStudy (anonymous):

and put into 2

OpenStudy (anonymous):

you should get kx+4y -2x-2y=0 4x+ky+2(-2x-2y)=0

OpenStudy (anonymous):

eq(1) kx+2y-2x=0 eq(2) ky-4y=0 y(k-4)=0 , y=0 or k=4

OpenStudy (anonymous):

(4)x+2y-2x=0 2x+2y=0 ----> x=-y

OpenStudy (anonymous):

Ahaa ! Thanks a lot ! :D This question was asked in the chapter vectors and 3-D , Do you think there is any other approach possible to this question ?

OpenStudy (anonymous):

@AJ01

OpenStudy (loser66):

when 3 planes intersects at a line, the rank of the augmented matrix must be 2.

OpenStudy (loser66):

To get it, among 3 equations, you must have one of them is the multiple of the other one. Try it

OpenStudy (loser66):

I believe that you have Linear Algebra course, right? If not, you don't have this question. It is about either linear algebra or calculus3 (dealing with planes)

OpenStudy (anonymous):

Umm , I do not really understand much of what you are saying . I do not have these course , but could you still please try to explain it to me ?

OpenStudy (loser66):

http://www.emathematics.net/posreltresplanos.php